Can you answer this for me
Elimination: 3x+2y=12
5x-4y=-2

Answer by Cromlix(4381)

3x + 2y = 12......1
5x - 4y = -2......2
Multiply (1) by 2
6x + 4y = 24......3
5x - 4y = -2......2
Add (1) to(2)
11x = 22
x = 2
Substitute x = 2
into Eq(1)
3x + 2y = 12
3(2) + 2y = 12
6 + 2y = 12
2y = 12 - 6
2y = 6
y = 3
x = 2 y = 3
Hope this helps.
:-)